OpenStudy (anonymous):

The length of a rectangle is 2 more than twice the width. Find the dimensions of the rectangle if the perimeter is 76. If x represents the width of the rectangle, then which expression represents the perimeter?

OpenStudy (anonymous):

these are my answer chooses 2x + 2 4x + 4 6x + 4

OpenStudy (texaschic101):

2x + 2

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

2x + 2 = the length

OpenStudy (anonymous):

so is it 4x+4 or 2x+2

OpenStudy (texaschic101):

x = width P = 2L + 2x L = 2x + 2 76 = 2(2x + 2) + 2x 76 = 4x + 4 + 2x 76 = 6x + 4 76 - 4 = 6x 72 = 6x x = 72/6 x = 12 (width) L = 2x + 2 L = 2(12) + 2 L = 24 + 2 L = 26 check... P = 2L + 2x 76 = 2(26) + 2(12) 76 = 52 + 24 76 = 76 (correct) width(x) = 12 Length (L) = 26

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

what one is is

OpenStudy (texaschic101):

Perimeter = 6x + 4

OpenStudy (anonymous):

ok thank you